1) Suppose that the elasticity of supply for wheat is 0.3 and for magazines is 1.3.

a. If the price of wheat rises 10%, what is the increase in quantity of wheat supplied?

b. If the price of of a magazine rises 10%, what is the increase in quantity of magazine supplied?

c. Which good does the quantity supplied respon more strongly, wheat or magazines?


1
Expert's answer
2022-01-26T10:22:32-0500

"\text{a.\, E.S}=\dfrac{\Delta Q_s}{\Delta P}\\
\text{substituting for E.S and Change in Price}\\ 
\text{we have;}\\
0.3=\dfrac{\Delta Q_s}{10}\\
\text{Cross multiplying, we have;}\\
\Delta Q_S=0.3*10\\
\Delta Q_S=3\\
\text{b.\, E.S}=\dfrac{\Delta Q_s}{\Delta P}\\
\text{Substituting for E.S and change in price,}\\
\text{we have;}\\
1.3=\dfrac{\Delta Q_s}{10}\\
\text{cross multiplying, we have;}\\
\Delta Q_S=1.3*10\\
\Delta Q_s=13\\
\text{c.\ Magazine responds more strongly to supply.}\\
\text{Magazine has a more elastic supply than Wheat.}\\
\text{Note: E.S means Elasticity of supply}\\
\Delta Q_s\ \text{means change in quantity supplied}\\
 \Delta P\ \text{means change in price}\\
\Delta P\ \text{means change in price}"